OpenStudy (anonymous):

n to the power of 3 = 15.625

OpenStudy (saifoo.khan):

\[n^3 = 15.625\]??

OpenStudy (anonymous):

yes

OpenStudy (saifoo.khan):

take a cube on both sides.

OpenStudy (saifoo.khan):

\[\sqrt[3]{n^3}=\sqrt[3]{15.625}\]

OpenStudy (anonymous):

\[\sqrt[3]{15.625} tht will get you 2.5\] stupid thing cube root

OpenStudy (saifoo.khan):

n = 2.5

OpenStudy (anonymous):

saifoo..khan is pro